OpenStudy (anonymous):

expand log 7 (3x-2)^2

OpenStudy (anonymous):

no. i figured it out

Nnesha (nnesha):

quotient rule\[\huge\rm log_b y - \log_b x = \log_b \frac{ x }{ y}\] to condense you can change subtraction to division product rule \[\huge\rm log_b x + \log_b y = \log_b( x \times y )\] addition ----> multiplication power rule \[\huge\rm log_b x^y = y \log_b x\] CHANGE OF BASE FORMULA: change of base formula \[\huge\rm log_a b = \frac{ \log b }{ \log a }\] anyways log properties if you need...